points a and b are 60 km. a part on a highway. a car starts from a and another car starts from b simultaneously. if they travel in the same direction ,they meet in 7 hours , but if they travel towards each other they meet in one hour. find the speed of two
Answer by lwsshak3(11628)

points a and b are 60 km. a part on a highway. a car starts from a and another car starts from b simultaneously. if they travel in the same direction ,they meet in 7 hours , but if they travel towards each other they meet in one hour. find the speed of two
***
let x=speed of faster car
let y=speed of slower car
travel time=distance/speed
..
60/(x+y)=1(travel toward each other)
60/(x-y)=7(travel in same direction)
..
x+y=60
7x-7y=60
..
7x+7y=420
7x-7y=60
add
14x=480
x=34.28
y=60-x=25.72
speed of faster car≈34.28 mph
speed of slower car≈25.72 mph